OpenStudy (anonymous):

get the y" of this please y = sqr. (1-4x)^3

OpenStudy (anonymous):

y = sqr. (1-4x)^3\[y = \sqrt{1-4x)^3}\]

OpenStudy (anonymous):

3/2(1-4x)^1/2 (-4)

OpenStudy (radar):

\[y ^{'}=-6\sqrt{1-4x}\] \[y ^{''}=12/\sqrt{1-4x}\]

OpenStudy (radar):

Based on:\[y=\sqrt{(1-4x)^{3}}=(1-4x)^{3/2}\]